A horizontal force 39ˍN is used to help push an 31ˍkg box up a 13ˍm incline with
angle 23°, as shown. What is the work done by this applied force? What is the work done by gravity?
1
Expert's answer
2012-11-19T10:41:22-0500
W = Delta PE = m*g* (delta h) = (31 кг) * (9.8 м / с / с) * (13 м) = 3 949,4 joule